OpenStudy (bunnykitten02):

Please help ! State whether either extreme in the data set is an outlier. A. The upper extreme is an outlier. B. Neither extreme is an outlier. C. Both extremes are outliers. D. The lower extreme is an outlier.

OpenStudy (bunnykitten02):

please help i have 3 more questions after this

OpenStudy (kropot72):

An outlier is an observation that is either (a) above the Upper Quartile + (1.5 * the Interquartile Range); or (b) below the Lower Quartile - (1.5 * the Interquartile Range).

OpenStudy (kropot72):

From the boxplot: UQ = 90 LQ = 68 IQR = UQ - LQ = 90 - 68 = 22 The upper cut-off is: \[\large UQ+1.5\times IQR=90+1.5\times22=90+33=123\] The lower cut-off is: \[\large LQ-1.5\times IQR=68-1.5\times22=35\]